OpenStudy (anonymous):

If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system?

OpenStudy (whpalmer4):

Let's look at an example: \[x+y = 1\]\[-x-y=-1\] If we add the two equations together, or use substitution, after the dust settles, we're left with \[0=0\] Now, observe that if we take the first equation and multiply it by -1, we get the second equation. What does that mean? Well, it means that the two lines are identical. A solution is found wherever the same point satisfies all the equations in a system, so any point on the line is a solution. There are infinitely many points on the line, so we have infinitely many solutions. If you solve both equations for \(y\), you'll see that you do have the same formula each time. Let's look at another very closely related example: \[x+y =1\]\[x+y=0\]Here if we solve each equation for \(y\), we get respectively\[y= -x+1\]\[y=-x\] If we think about slope-intercept form \((y=mx+b)\), both lines have the same slope, but a different y-intercept. That means they are parallel, and parallel lines do not intersect. What happens if we solve the system? I'll solve the system by plugging in the expression for \(y\) from the second solution in the first equation: \[x + y = 1\]\[x + (-x) = 1\]\[0 = 1\] When we get a result like this, it means the lines have the identical slope but have different y-intercepts (parallel), distinguishing from the case of lines with the identical slope and identical y-intercepts (coincident). To sum up, there are three cases: 1) 0 = 0 Lines are coincident, infinitely many solutions 2) 0 = 1 (or any number other than 0) Lines are parallel, no solutions 3) (x,y) (for some value of x and y) Lines intersect at a point; that point is a solution to the system